# Software Fairness Testing in Practice

### v1

@@ Abstract
Software testing ensures that a system functions correctly, meets specified requirements, and maintains high quality. As artificial intelligence and machine learning (ML) technologies become integral to software systems, testing has evolved to address their unique complexities. A critical advancement in this space is fairness testing, which identifies and mitigates biases in AI applications to promote ethical and equitable outcomes. Despite extensive academic research on fairness testing—including test input generation, test oracle identification, and component testing—practical adoption remains limited. Industry practitioners often lack clear guidelines and effective tools to integrate fairness testing into real-world AI development. This study investigates how software professionals test AI-powered systems for fairness through interviews with 22 practitioners working on AI and ML projects. Our findings highlight a significant gap between theoretical fairness concepts and industry practice. While fairness definitions continue to evolve, they remain difficult for practitioners to interpret and apply. The absence of industry-aligned fairness testing tools further complicates adoption, necessitating research into practical, accessible solutions. Key challenges include data quality and diversity, time constraints, defining effective metrics, and ensuring model interoperability. These insights emphasize the need to bridge academic advancements with actionable strategies and tools, enabling practitioners to systematically address fairness in AI systems.
@@ I Introduction
In the current landscape of software engineering, fairness testing has emerged as a crucial component of quality assurance, addressing the ethical challenges posed by AI-driven systems. These systems, capable of performing complex tasks such as recommendation, object detection, classification, and prediction, require advanced validation techniques to ensure their reliability and fairness [6, 7]. Unlike traditional software, AI models continuously adapt to data, making them susceptible to biases introduced through training data, algorithmic design, or societal inequalities [8]. Fairness testing, then, aims to systematically evaluate the ethical implications and potential biases embedded within AI systems, ensuring that decision-making processes lead to equitable outcomes across diverse demographic groups. By identifying and mitigating discriminatory risks based on factors such as race, gender, or socioeconomic status, fairness testing plays a key role in ensuring ethical AI development and promoting fairness in automated decision-making [9, 10].
While extensive research has been conducted on fairness testing, focusing on areas such as test input generation, test oracle identification, and the testing of various components—including data testing, machine learning program testing, and model testing [11]—much of this work remains confined to academic settings, offering limited practical guidance for industry professionals. This gap leaves practitioners struggling to effectively implement fairness testing and manage fairness-related challenges in real-world AI systems. For example, there is a growing demand for fairness testing tools that align more closely with industry needs, particularly in terms of usability, comprehensive coverage, and configurability, ensuring they can be seamlessly integrated into existing development workflows [12].
@@ II-A Software Fairness
Although there is no single, universally accepted definition of fairness in the software context, the term generally refers to a set of properties designed to prevent discriminatory behaviors or biases, particularly regarding sensitive attributes such as gender, race, and age [13, 14, 7]. Ensuring fairness is essential for all types of software systems, especially AI-driven applications used in critical domains like criminal justice, healthcare, and recruitment [6]. Without proper fairness considerations, these systems can perpetuate or even amplify existing social inequalities, leading to biased and discriminatory outcomes that disproportionately impact certain individuals or groups [14].
@@ II-B Fairness Testing
Despite efforts to design discrimination-free algorithms, bias often persists due to implementation errors, emergent system behaviors, and automated learning from data [9]. This makes fairness testing essential for identifying and mitigating system bias [16]. Fairness Testing is the process of verifying and validating the software to reveal fairness bugs, which are flaws that cause discrepancies between actual behavior and fairness conditions [11]. In other words, it ensures that software operates impartially, without discriminating against individuals or groups based on sensitive attributes such as sex, race, or age [11].
@@ III-A The Case
Using structured interviews as our main data collection approach, we interviewed a group of software professionals directly involved in the development of AI and machine learning-based software solutions, including designers, software engineers, data scientists, and testers, from four different projects: Project A developed a real-time application that translates sign language into Portuguese text using deep learning neural networks [24]. Project B involved creating digital twins and predictive models for the petroleum and energy sector, Project C utilized large language models (LLMs) in education, and Project D employed computer vision for facial recognition [25]). While specific details about Projects B and C are restricted due to confidentiality agreements, all projects, regardless of their individual or industrial focus, share a reliance on AI technologies. This reliance highlights the need to address bias issues and ensure fairness, or bias testing is implemented to avoid discriminatory or biased outcomes and ensure trustworthy AI systems.
@@ III-C Data Collection
As part of our data collection process, we used a semi-structured interview guide (Table I) that evolved across the three interview rounds. In the initial stage, conversations focused on general concepts of bias in AI development. As new themes emerged, subsequent rounds incorporated more targeted questions about fairness testing, such as challenges related to integrating fairness considerations into development cycles, methods used to detect bias, the roles responsible for fairness validation, and how practitioners address fairness concerns throughout the software lifecycle. This iterative refinement enabled us to capture a deeper understanding of how fairness testing is operationalized in practice and the obstacles professionals encounter in real-world projects.
@@ III-E Data Analysis
We followed a three-phase thematic analysis approach: (1) Open coding, where GPT-4 Omni was prompted to extract fairness testing references across transcripts; (2) Axial coding, where we manually grouped codes into broader categories based on semantic similarity and development context; and (3) Selective coding, where we synthesized these categories into broad themes.
We began with open coding, examining transcripts to identify AI testing practices and generate initial codes that described different aspects of the data. To improve efficiency in this phase, we applied GPT-4 Omni, following guidelines for using LLMs in qualitative analysis [29, 30]. We used prompt engineering (Table II) to extract relevant quotations from the interviews, iteratively refining prompts and manually verifying results to ensure the model accurately identified and categorized fairness testing-related mentions [31, 32, 33]. The prompt shown in Table II reflects the final version, refined through multiple iterations. Manual revision was performed throughout this phase to iteratively refine or reword GPT-generated codes. Once consistency was achieved across a test set of transcripts, we finalized the prompt and proceeded with full-scale extraction. To validate the results, we randomly selected 20% of the retrieved quotes and manually verified them against the original transcripts, confirming that the extracted content accurately reflected participant responses.
@@ IV Findings
The final demographic breakdown of our sample reflects this diversity. Approximately 32% of participants were non-male professionals, including one non-binary individual. Additionally, 8% of participants had disabilities, with two individuals reporting a disability. The sample also included 16% non-white professionals, with representation from mixed-race and Black individuals. 20% identified as LGBTQIA+, comprising gay, lesbian, and queer participants. Furthermore, 12% were neurodivergent, including individuals with ADHD. In terms of professional expertise, 40% held advanced degrees (Master’s or Ph.D.), and 40% had over five years of experience in software development. This diverse and highly skilled group provided valuable perspectives on how diversity influences AI development and helps address bias. Further details are provided in Table III.
@@
Education Bachelor’s Degree 8 individuals
Postbaccalaureate 3 individuals
@@ IV-A Fairness Testing in Practice: What is Being Tested
Based on our analysis, we identified three distinct targets for fairness testing that practitioners actively develop to detect biases and mitigate fairness issues. The first target, testing for diversity and inclusivity, focuses on ensuring software performs fairly across diverse demographics, physical attributes, and cultural contexts. Practitioners in this category conduct detailed testing that includes sensitive attributes such as ethnicity, gender, and socio-economic background, validating that AI systems treat all demographic groups equitably. This user-centered approach highlights the social implications of fairness testing, aligning technical objectives with broader ethical considerations.
The second focus, testing for model consistency, involves verifying the AI model’s robustness, particularly when handling unfamiliar or unexpected inputs. Practitioners aim to ensure the model generalizes beyond its training data, producing stable and unbiased outputs across a range of conditions. This approach addresses the need for AI systems to function reliably in real-world scenarios, reducing the risk of inconsistent or biased responses. Professionals engaged in this testing strategy experiment with varied inputs, including edge cases, to validate model behavior across diverse situations.
Finally, testing for data representation and balancing centers on ensuring that training data is representative and equitably distributed across user groups. Practitioners work to prevent bias at the data level by identifying overrepresentation or underrepresentation in key demographic categories. This approach is critical for mitigating embedded biases that might otherwise go undetected, helping ensure that models do not disproportionately favor or disadvantage any particular group.
@@
For example, one participant stated, “We did not thoroughly test how the model would react to extreme changes in data distribution, such as large lighting variations” (P17), while another noted, “We don’t examine the impact of extreme changes in data, such as concept drift, in all cases” (P18). Others raised similar concerns, indicating that “we did not investigate the effects of significant environmental variations, which may have impacted results” (P15, P19). These findings highlight an area in need of attention to ensure fairness testing accounts for real-world variability.
@@ IV-B Fairness Testing in Practice: Strategies and Industry Approaches
The second approach, manipulation and simulation of data, broadens test inputs to reflect real-world variations and capture edge cases. Manipulation involves altering data characteristics—such as changing user attributes or environmental conditions—to assess their impact on model performance. Simulation, on the other hand, generates synthetic data to represent conditions not present in the original dataset. Both techniques help enrich test coverage and completeness, often guided by testing professionals’ expertise.
The third approach relies on well-defined metrics to systematically assess fairness and detect bias. Practitioners use metrics such as accuracy, recall, and precision to evaluate model outputs, ensuring alignment with fairness objectives. Additionally, A/B testing is employed to compare model behaviors under varied conditions, helping practitioners identify biases affecting specific user groups. In practice, data scientists and programmers manage most of these metrics, while testing professionals primarily apply A/B testing to assess fairness.
@@
Despite these emerging strategies, fairness testing in practice remains heavily reliant on verification processes conducted by data scientists and programmers, with limited involvement from dedicated testing professionals. As a result, critical decisions about fairness testing are often made without fully integrating specialized testing insights, leading to a narrower perspective on fairness-related issues. Additionally, industry-driven practices and tools specifically designed for fairness testing are still lacking, leaving many testing efforts inconsistent and ad-hoc.
@@ V-A Software Fairness Testing: The Gap Between Academic Research and Practice
While researchers have made significant contributions in defining individual and group fairness—introducing concepts such as fairness through unawareness, awareness, counterfactual fairness, and causal fairness for individual fairness, and statistical parity, equalized odds, and equal opportunity for group fairness [11, 14, 35]—these advancements are yet to be fully integrated into industry practice. In theory, these definitions offer structured approaches for ensuring equitable outcomes across different demographic groups and for making fair individual comparisons, which could ideally serve as the foundation for standardized testing processes. However, our findings reveal that, in practice, professionals still rely on ad-hoc methods to implement fairness testing due to a lack of clear, actionable guidelines. This gap between theoretical frameworks and real-world applications underscores the urgent need for practical, standardized tools and guidelines to make fairness definitions operational within the software industry.
@@
The literature highlights a growing number of open-source fairness testing tools available for ML applications, with several tools identified across domains like general ML, deep learning, natural language processing, computer vision, and speech recognition [11]. However, key limitations—such as steep learning curves, insufficient documentation, limited configurability, and lack of regular maintenance—hinder their adoption in real-world industry settings [12]. Our findings echo these issues, revealing that, despite the documented tools in literature, these resources are rarely incorporated into everyday fairness testing practices in the software industry. This gap between tool availability and practical usage demonstrated the need for tools that are more aligned with industry demands and practitioner needs. In particular, given the knowledge gaps around fairness testing methods reported by practitioners, tools designed for ease of use could play a significant role in making fairness testing more accessible and actionable within industry practices.
@@ V-B Implications for Research
While our findings confirm the persistent gap between fairness theory and its practical implementation, they also point to actionable research directions to bridge this divide. Theoretical progress in defining fairness has introduced structured approaches such as fairness through awareness, counterfactual fairness, and statistical parity. However, these definitions remain challenging for practitioners to apply due to their complexity and lack of direct integration into existing software development workflows. To address this, research should prioritize the development of frameworks that translate fairness definitions into actionable methodologies that software teams can incorporate seamlessly into testing processes.
Moreover, our study highlights the need for comprehensive, industry-aligned fairness testing tools that go beyond academic prototypes. While the literature has produced numerous tools for bias detection and fairness validation, their adoption in industry remains limited due to high technical barriers, lack of configurability, and insufficient documentation. Research should focus on designing user-friendly, well-documented, and scalable tools that integrate with industry-standard workflows. Additionally, empirical studies assessing the effectiveness of fairness testing methodologies in real-world settings could provide valuable insights into best practices and gaps that require further attention.
@@ V-C Implications for Practice
Another key consideration for industry is the integration of fairness testing tools into standard testing pipelines. As our findings indicate, fairness testing is currently performed predominantly by data scientists and programmers, with limited involvement from dedicated testing professionals. To ensure fairness testing becomes a sustainable practice, organizations should adopt tools that enable seamless integration into existing automated testing environments. Additionally, fostering collaboration between different roles—such as software testers, data scientists, and domain experts—can enhance the effectiveness of fairness assessments and lead to more robust, bias-aware AI systems.

### v2

@@ Abstract
Software testing ensures that a system functions correctly, meets specified requirements, and maintains high quality. As artificial intelligence and machine learning technologies become integral to software systems, testing has evolved to address their unique complexities. A critical advancement in this space is fairness testing, which identifies and mitigates biases in AI applications to promote ethical and equitable outcomes. Despite extensive academic research on fairness testing, including test input generation, test oracle identification, and component testing, practical adoption remains limited. Industry practitioners often lack clear guidelines and effective tools to integrate fairness testing into real-world AI development. This study investigates how software professionals test AI-powered systems for fairness through interviews with 22 practitioners working on AI and ML projects. Our findings highlight a significant gap between theoretical fairness concepts and industry practice. While fairness definitions continue to evolve, they remain difficult for practitioners to interpret and apply. The absence of industry-aligned fairness testing tools further complicates adoption, necessitating research into practical, accessible solutions. Key challenges include data quality and diversity, time constraints, defining effective metrics, and ensuring model interoperability. These insights emphasize the need to bridge academic advancements with actionable strategies and tools, enabling practitioners to systematically address fairness in AI systems.
@@ I Introduction
In the current landscape of software engineering, fairness testing has emerged as a key component of quality assurance, addressing the ethical challenges posed by AI-driven systems. These systems, capable of performing complex tasks such as recommendation, object detection, classification, and prediction, require advanced validation techniques to ensure their reliability and fairness [6, 7]. Unlike traditional software, AI models continuously adapt to data, making them susceptible to biases introduced through training data, algorithmic design, or societal inequalities [8]. Fairness testing, then, aims to systematically evaluate the ethical implications and potential biases embedded within AI systems, ensuring that decision-making processes lead to equitable outcomes across diverse demographic groups. By identifying and mitigating discriminatory risks based on factors such as race, gender, or socioeconomic status, fairness testing plays a key role in ensuring ethical AI development and promoting fairness in automated decision-making [9, 10].
While extensive research has been conducted on fairness testing, focusing on areas such as test input generation, test oracle identification, and the testing of various components, including data testing, machine learning program testing, and model testing [11], much of this work remains confined to academic settings, offering limited practical guidance for industry professionals. This gap leaves practitioners struggling to effectively implement fairness testing and manage fairness-related bugs in real-world AI systems [9]. For example, there is a growing demand for fairness testing tools that align more closely with industry needs, particularly in terms of usability, comprehensive coverage, and configurability, ensuring they can be seamlessly integrated into existing development workflows [12].
@@ II-A Software Fairness
Although there is no single, universally accepted definition of fairness in the software context, the term generally refers to a set of properties designed to prevent discriminatory behaviors or biases, particularly regarding sensitive attributes such as gender, race, and age [13, 14, 7]. Ensuring fairness is essential for all types of software systems, especially AI-driven applications used in critical domains like criminal justice, healthcare, and finances [6]. Without proper fairness considerations, these systems can perpetuate or even amplify existing social inequalities, leading to biased and discriminatory outcomes that disproportionately impact certain individuals or groups [14].
@@ II-B Fairness Testing
Despite efforts to design discrimination-free algorithms, bias often persists due to implementation errors, emergent system behaviors, and automated learning from data [9]. This makes fairness testing essential for identifying and mitigating system bias [16]. Fairness Testing is the process of verifying and validating the software to reveal fairness bugs, which are flaws that cause discrepancies between actual behavior and fairness conditions [11]. In other words, it ensures that software operates impartially, without discriminating against individuals or groups based on sensitive attributes such as gender, race, or age [11].
@@ III-A The Case
Using semi-structured interviews as our main data collection approach, we interviewed a group of software professionals directly involved in the development of AI and machine learning-based software solutions, including designers, software engineers, data scientists, and testers, from four different projects: Project A developed a real-time application that translates sign language into Portuguese text using deep learning neural networks [24]. Project B involved creating digital twins and predictive models for the petroleum and energy sector, Project C utilized large language models (LLMs) in education, and Project D employed computer vision for facial recognition [25]). While specific details about Projects B and C are restricted due to confidentiality agreements, all projects, regardless of their individual or industrial focus, share a reliance on AI technologies. This reliance highlights the need to address bias issues and ensure fairness, or bias testing is implemented to avoid discriminatory or biased outcomes and ensure trustworthy AI systems.
@@ III-C Data Collection
As part of our data collection process, we used a semi-structured interview guide (Table I) that evolved across three interview rounds. In the initial stage, conversations focused on general concepts of bias in AI development. As new themes emerged, subsequent rounds incorporated more targeted questions about fairness testing, such as challenges related to integrating fairness considerations into development cycles, methods used to detect bias, the roles responsible for fairness validation, and how practitioners address fairness concerns throughout the software lifecycle. This iterative refinement enabled us to capture a deeper understanding of how fairness testing is operationalized in practice and the obstacles professionals encounter in real-world projects.
@@ III-E Data Analysis
We followed a three-phase thematic analysis approach as demonstrated in Figure 1: (1) Open coding, where GPT-4 Omni was prompted to extract fairness testing references across transcripts; (2) Axial coding, where we manually grouped codes into broader categories based on semantic similarity and development context; and (3) Selective coding, where we synthesized these categories into broad themes.
We began with open coding, examining transcripts to identify AI testing practices and generate initial codes that described different aspects of the data. To improve efficiency in this phase, we applied GPT-4 Omni, following guidelines for using LLMs in qualitative analysis [29, 30]. Leveraging an LLM aimed to significantly accelerate the coding process, acknowledging that fully manual thematic analysis is typically time-consuming [31]. We used prompt engineering (Table II) to systematically extract pertinent quotations from interview transcripts. This process involved refining the prompt and manually verifying the results to ensure accurate identification and categorization of mentions related to fairness testing [32, 33, 34]. The prompt presented in Table II represents the finalized version, optimized by applying multiple prompt engineering techniques [30]. Manual review was performed to verify GPT-generated codes. After achieving consistency on our dataset, we conducted full-scale extraction. For validation, 20% of extracted quotes were randomly selected and manually cross-checked against original transcripts to confirm accuracy.
@@ IV Findings
The final demographic breakdown of our sample reflects this diversity. Approximately 32% of participants were non-male professionals, including one non-binary individual. Additionally, 8% of participants had disabilities, with two individuals reporting a disability. The sample also included 16% non-white professionals, with representation from mixed-race and Black individuals. 20% identified as LGBTQIA+, comprising gay, lesbian, and queer participants. Furthermore, 12% were neurodivergent, including individuals with ADHD. In terms of professional expertise, 36% held advanced degrees (Master’s or Ph.D.), and 40% had more than five years of experience in software development. This diverse and highly skilled group provided valuable perspectives on how diversity influences AI development and helps address bias. Further details are provided in Table III.
@@
Education Bachelor’s Degree 9 individuals
Postbaccalaureate 5 individuals
@@ IV-A Fairness Testing in Practice: What is Being Tested
Based on our analysis, we identified three distinct targets for fairness testing that practitioners actively use to detect biases and mitigate fairness issues. The first target, testing for diversity and inclusivity, focuses on ensuring software performs fairly across diverse demographics, physical attributes, and cultural contexts. Practitioners discussing this target conduct detailed testing that includes sensitive attributes such as ethnicity, gender, and socio-economic background, validating that AI systems treat all demographic groups equitably. This user-centered approach highlights the social implications of fairness testing, aligning technical objectives with broader ethical considerations.
The second focus, testing for model consistency, involves verifying the AI model’s robustness, particularly when handling unfamiliar or unexpected inputs. Looking at this target, practitioners aim to ensure the model generalizes beyond its training data, producing stable and unbiased outputs across a range of conditions. This approach addresses the need for AI systems to function reliably in real-world scenarios, reducing the risk of inconsistent or biased responses. Professionals engaged in this testing strategy experiment with varied inputs, including edge cases, to validate model behavior across diverse situations.
Finally, testing for data representation and balancing centers on ensuring that training data is representative and equitably distributed across user groups. Practitioners focusing on this target work to prevent bias at the data level by identifying overrepresentation or underrepresentation in key demographic categories. This approach is necessary for mitigating embedded biases that might otherwise go undetected, helping ensure that models do not disproportionately favor or disadvantage any particular group.
@@
For example, one participant stated, “We did not thoroughly test how the model would react to extreme changes in data distribution, such as large lighting variations” (P17), while another noted, “We don’t examine the impact of extreme changes in data, such as concept drift, in all cases” (P18). Others raised similar concerns, indicating that they did not investigate the effects of significant environmental variations, which may have impacted testing results (P15, P19). These findings highlight an area in need of attention to ensure fairness testing accounts for real-world variability.
@@ IV-B Fairness Testing in Practice: Strategies and Industry Approaches
The second approach, manipulation and simulation of data, broadens test inputs to reflect real-world variations and capture edge cases. Manipulation involves altering data characteristics, such as changing user attributes or environmental conditions, to assess their impact on model performance. Simulation, on the other hand, generates synthetic data to represent conditions not present in the original dataset. Both techniques help enrich test coverage and completeness, often guided by testing professionals’ expertise.
The third approach utilizes well-defined metrics to systematically evaluate fairness and identify bias. Data scientists and programmers apply metrics such as accuracy, recall, and precision to assess model outputs and ensure alignment with fairness objectives. Additionally, A/B testing is conducted to compare model behavior under varying conditions, enabling testers to identify biases affecting specific user groups within controlled experimental settings.
@@
Despite these emerging strategies, fairness testing in practice remains heavily reliant on verification processes conducted by data scientists and programmers, with limited involvement from dedicated testing professionals. This limited involvement occurs primarily because existing methods focus mainly on dataset design and model evaluation. As a result, critical decisions about fairness testing are often made without fully integrating specialized testing insights, leading to a narrower perspective on fairness-related issues. Additionally, industry-driven practices and tools specifically designed for fairness testing are still lacking, leaving many testing efforts inconsistent and ad-hoc.
@@ V-A Software Fairness Testing: The Gap Between Academic Research and Practice
While researchers have made significant contributions in defining individual and group fairness by introducing concepts such as fairness through unawareness, awareness, counterfactual fairness, and causal fairness for individual fairness, and statistical parity, equalized odds, and equal opportunity for group fairness [11, 14, 36], these advancements are yet to be fully integrated into industry practice. In theory, these definitions offer structured approaches for ensuring equitable outcomes across different demographic groups and for making fair individual comparisons, which could ideally serve as the foundation for standardized testing processes. However, our findings reveal that, in practice, professionals still rely on ad-hoc methods to implement fairness testing due to a lack of clear, actionable guidelines. This gap between theoretical frameworks and real-world applications highlights the urgent need for practical, standardized tools and guidelines to operationalize fairness definitions within the software industry.
@@
The literature shows a growing number of open-source fairness testing tools available for ML applications, with several tools identified across domains like general ML, deep learning, natural language processing, computer vision, and speech recognition [11]. However, limitations such as steep learning curves, insufficient documentation, limited configurability, and lack of regular maintenance hinder their adoption in real-world industry settings [12]. Our findings echo these issues, revealing that, despite the documented tools in literature, these resources are rarely incorporated into everyday fairness testing practices in the software industry. This gap between tool availability and practical usage demonstrated the need for tools that are more aligned with industry demands and practitioner needs. In particular, given the knowledge gaps around fairness testing methods reported by practitioners, tools designed for ease of use could play a significant role in making fairness testing more accessible and actionable within industry practices.
@@ V-B Implications for Research
While our findings confirm the persistent gap between fairness theory and its practical implementation, they also point to several research directions to bridge this divide. Theoretical progress in defining fairness has introduced structured approaches such as fairness through awareness, counterfactual fairness, and statistical parity. However, these definitions remain challenging for practitioners to apply due to their complexity and lack of direct integration into existing software development workflows. To address this, research should prioritize the development of frameworks that translate fairness definitions into actionable methodologies that software teams can incorporate seamlessly into testing processes.
Moreover, our study highlights the need for comprehensive, industry-aligned fairness testing tools that go beyond academic prototypes. While the literature has produced numerous tools for bias detection and fairness validation, their adoption in industry remains limited due to high technical barriers, lack of configurability, and insufficient documentation. Research should focus on designing user-friendly, well-documented, and scalable tools that integrate with industry-standard workflows. Additionally, empirical studies assessing the effectiveness of fairness testing methodologies in real-world settings could provide valuable results into best practices and gaps that require further attention.
@@ V-C Implications for Practice
Another key consideration for industry is the integration of fairness testing tools into standard testing pipelines. As our findings indicate, fairness testing is currently performed predominantly by data scientists and programmers, with limited involvement from dedicated testing professionals. To ensure fairness testing becomes a sustainable practice, organizations should adopt tools that enable seamless integration into existing automated testing environments. Additionally, fostering collaboration between different roles (e.g., software testers, data scientists, and domain experts) can enhance the effectiveness of fairness assessments and lead to more robust, bias-aware AI systems.